OpenStudy (anonymous):

The cost of a cell phone bill (C) increases when the number of text messages (T) increases. Write the correct equation for this scenario and solve for the cost when the number of texts is 4. C = 10 over t; C = 20 C = 10 over t; C = 5 C = .8T; C = 16 C = .4T; C= 1.6

OpenStudy (anonymous):



OpenStudy (anonymous):

@austinL can you help with this?

OpenStudy (anonymous):

@phi i need help understanding this ^^ please

OpenStudy (anonymous):

@eninone i need understanding this please... btw you are an awesome helper

OpenStudy (anonymous):

@nerlineg you've a curiosity to learn, that's the power of learning. btw, i'm just a beginner in math, there's still to learn/understand much much the mighty universe of math.

OpenStudy (anonymous):

\[as \ the\ cost \ of \ telephone\ bill\ (C)\ increases\\ when\ the\ number\ of\ texts \ messages (T) \ increases.\\ \implies\ C \ \alpha \ T\ [ C \ is\ directly\ proportional \ with\ T]\\ \implies\ C=kT,\ where\ k\ is\ constant\\ find\ k \ \it\ is\ equal \ 5/2\\ so\ C=\frac{ 5 }{ 2}T---(1)\\ a/c \ problem \ requirement\ when\ T=4\\then\ from\ eq(1),\\ C=\frac{ 5 }{ 2 }\times4\\implies\ C=10\]

OpenStudy (anonymous):